Please show me how to get the square root. Thanks
(5 square root of 6-4)(3square root of 6-1)

Answer by stanbon(75887)

(5 square root of 6-4)(3square root of 6-1)
--------
(5sqrt6 -4 )(3sqrt6 -1)
Use FOIL to get this product:
F: (5sqrt6*3sqrt6)= 15*6=90
O: (5sqrt6 *-1) = -5sqrt6
I: (-4*3sqrt6) = -4sqrt6
L: -4*-1 = 4
---------------
Add the "like" terms to get
= 94 - 9sqrt6
===========
Cheers,
Stan H.